### v1

# Searching for axion dark matter with magnetic resonance force microscopy

Abstract
We propose a magnetic resonance force microscopy (MRFM) search for axion dark matter around 1 $\rm GHz$. The experiment leverages the axion’s derivative coupling to electrons, which induces an effective A.C. magnetic field on a sample of electron spins polarized by a D.C. magnetic field and a micromagnet. A second pump field at a nearby frequency enhances the signal, with the detuning matched to the resonant frequency of a magnet-loaded mechanical oscillator. The resulting spin-dependent force is detected with high sensitivity via optical interferometry. Accounting for the relevant noise sources, we show that current technology can be used to put constraints competitive with those from laboratory experiments with just a minute of integration time. Furthermore, varying the pump field frequency and D.C. magnetic field allows one to scan the axion mass. Finally, we explore this setup’s capability to put constraints on other dark matter - Standard Model couplings.
[…]
While dark matter is known to interact with baryonic matter gravitationally, ultra-weak non-gravitational couplings are present in most theoretical models. Many of these couplings have effects that manifest as a real or effective A.C. magnetic field oscillating at $\omega_{\rm DM}$. For instance, the axion-photon coupling and the dark photon induce effective currents that generate real electromagnetic fields sikivie1983experimental; wilczek1987two; arza2022earth; chaudhuri2015radio. Furthermore, the axion-fermion couplings induce effective magnetic fields on fermion spins barbieri1989axion; alexander2018detecting. Due to the wide viable mass range for ULDM around 20 orders of magnitude, various technologies are required to fully search the allowed parameter space. This underscores a critical need for the development of ultra-precise A.C. magnetometers across a vast frequency range for dark matter detection. Several dark matter searches are based on magnetic resonance techniques in magnetized samples crescini2018operation; crescini2020axion; flower2019broadening; ikeda2022axion; jackson2020overview; garcon2019constraints; walter2025search; xu2024constraining. Various other magnetometry techniques have been used to search for dark matter bloch2020axion; bloch2023constraints; lee2023laboratory; wei2025dark; bloch2022new; abel2023search; gao2022axion; karanth2023first; jiang2021search; gavilan2024searching. In this manuscript, we propose a new mechanical, frequency-tunable search for dark matter using magnetic resonance force microscopy (MRFM).
Figure 1: Diagram of the MRFM setup. The top illustrates the general scenario where a magnetized sample under the influence of a D.C. field $B_{0}\hat{z}$ and a field due to a nearby micromagnet attached to a mechanical resonator undergoes magnetic resonance electron spins in the presence of a dark matter induced magnetic field $B_{\rm DM}\cos(\omega_{\rm DM}t)$. All the spins within the resonance region put a force on the resonator. A pump field $B_{\rm p}\cos(\omega_{\rm p}t)$ parallel to the $B_{\rm DM}$ amplifies the dark matter signal and modulates it down to the micromagnet-loaded resonator’s resonance frequency $|\omega_{\rm DM}-\omega_{\rm p}|=\omega_{\rm m}$, inducing vibrations that are read out optically. The bottom shows the two mechanical systems considered in this work: membranes and cantilevers.
We first recast various dark matter searches via various couplings into an A.C. magnetic field detection problem. Then, we introduce a magnetic resonance force microscopy sidles1991noninductive; sidles1995magnetic; suter2004magnetic detection scheme involving frequency down-conversion and optimize it for the axion-electron coupling. We show that the axion-electron coupling can be probed beyond current laboratory constraints with existing technology in under a minute of integration time. We then discuss the frequency scanning capabilities of our proposed setup. Finally, we explore this setup’s capability to put constraints on other dark matter - Standard Model couplings.
Dark Matter Induced Magnetic Fields.– Here we describe various ULDM couplings to Standard Model fields that manifest as effective or real magnetic field effects. Starting with axions, which are pseudoscalar particles that can be described by a classical wave of amplitude $\sqrt{2\rho_{\rm DM}}/\omega_{\rm DM}$ oscillating at its Compton frequency $\omega_{\rm DM}$. The axion field $a(\vec{x},t)$ can interact with fermions $\psi_{f}$ via the operator $g_{af}\partial_{\mu}a\bar{\psi}_{f}\gamma^{\mu}\gamma_{5}\psi_{f}$, where $g_{af}$ is the axion-fermion coupling strength, $\gamma^{\mu}$ are the Dirac matrices and $f=e,p,n$ for electrons, protons and neutrons. This leads to an interaction Hamiltonian of the form $\gamma_{f}\vec{S}\cdot\vec{B}_{\rm eff}$, where $\gamma_{f}=g_{f}e/2m_{f}$ is the fermion gyromagnetic ratio with the Lande g gactor $g_{f}$, $\vec{S}=\frac{\hbar}{2}\vec{\sigma}$ with Pauli matrices $\vec{\sigma}$, and $\vec{B}_{\rm eff}$ is an induced effective magnetic field barbieri1989axion; alexander2018detecting. For axion couplings to electrons, protons and neutrons, respectively, the effective magnetic field strengths $B_{af}$ are given below in terms of the coupling strength $g_{af}$:
$B_{ae}\approx 2\times 10^{-16}~\text{T}\times\frac{g_{ae}}{7\times 10^{-9}},$ (1)
$B_{ap}\approx 5\times 10^{-12}~\text{T}\times\frac{g_{ap}}{5\times 10^{-4}},$ (2)
$B_{an}\approx 1\times 10^{-11}~\text{T}\times\frac{g_{an}}{5\times 10^{-4}}.$ (3)
These equations give the magnetic field sensitivity required to probe the corresponding couplings to the level of current laboratory constraints in the $\rm GHz-THz$ range.
Axions can also couple to photons and induce an effective current that generates real magnetic fields sikivie1983experimental; wilczek1987two; caputo2021dark; arza2022earth. Here, a background magnetic field $B_{0}$ is required to mediate the interaction, and the induced magnetic field as a function of coupling strength $g_{a\gamma}$ is
$B_{a\gamma}=3\times 10^{-20}~\text{T}\times\left(\frac{L}{10~\text{cm}}\right)\left(\frac{B_{0}}{25~\rm\mu T}\right)\frac{g_{a\gamma}}{10^{-9}~\text{GeV}^{-1}},$ (4)
where $L$ is the size of the experiment. It is noteworthy that the axion-induced magnetic fields given above are independent of dark matter mass/frequency.
In addition to axions, dark photons coupling to Standard Model photons can also induce an effective current that generates real electromagnetic fields caputo2021dark; arza2022earth. The magnetic field generated by the dark photon as a function of the kinetic mixing parameter $\epsilon$ and experiment size $L$ is
$B_{A^{\prime}}=4\times 10^{-18}~\text{T}\times\left(\frac{f_{\rm DM}}{10^{9}~\rm Hz}\right)\left(\frac{L}{15~\rm cm}\right)\frac{\epsilon}{10^{-13}}.$ (5)
Unlike the axion couplings, the induced magnetic field here is dependent on dark matter mass/frequency. As before, these magnetic field strengths are written to show the required magnetometer sensitivity to match current laboratory constraints. While better astrophysical constraints may exist for certain couplings, these may be more model dependent, necessitating new laboratory searches to supplement these constraints. It is also worth reiterating that these equations give strengths of dark matter-induced magnetic fields $B_{\rm DM}$ which oscillate at $\omega_{\rm DM}$. See Appendix A for further details of the various dark matter induced magnetic fields. We now show how these weak signals can be detected by re-purposing an existing technology.
Magnetic Resonance Force Microscopy.– Magnetic Resonance Force Microscopy is a scanned-probe technique that combines imaging capabilities of magnetic resonance imaging with resolution of atomic force microscopy hammel2007magnetic. Unlike conventional magnetic resonance detection, which inductively detects the electromagnetic signal of transverse magnetization, MRFM directly senses changes in longitudinal magnetization through a mechanical resonator sidles1993signal. MRFM experiments generally employ cyclical modulation of longitudinal magnetization $M_{z}$ at the mechanical resonance frequency, using techniques like amplitude modulation (AM) Suter2004; wago1997magnetic. An advantage of this mechanical readout is avoiding circuit noise and associated backaction damping. Appendix B provides a more detailed comparison of MRFM with other magnetic resonance imaging techniques.
A typical setup involves a micromagnet attached to the mechanical resonator placed directly above the sample, as shown in Fig. 1. For an extended sample that is placed at a large enough distance from the micromagnet, the micromagnet can be treated as a point dipole suter2002probe; sidles91, as detailed in Appendix B. The micromagnet produces a strong local magnetic field gradient that couples the resonator to the spins of the magnetized sample underneath it. As a result of this interaction, the resonator senses a force given by Suter2004,
$F_{z}(t)=\int{-\left[\vec{m}(\vec{r},t)\cdot\vec{\nabla}\right]B_{z}(\vec{r})d^{3}r},$ (6)
with $\vec{m}(\vec{r},t)$ and $B_{z}(\vec{r})$ as magnetization density of the sample and the magnetic field of the micromagnet along the direction of mechanical motion, respectively. This integral is evaluated over the resonance area of the magnetized sample, as discussed below.
To induce magnetic spin resonance, a static magnetic field $B_{0}$ along the $z$ direction polarizes the sample, and an oscillating magnetic field $B_{\rm rf}$ is applied perpendicular to it. The bias field $B_{0}$ and the micromagnet’s magnetic field $B_{z}$ determine the Larmor frequency $\omega_{L}=\gamma(B_{0}+B_{z})$. Here, the role of the radio frequency field is played by the dark matter-induced magnetic field $B_{\rm DM}\cos(\omega_{\rm DM}t)$. To enhance the weak dark matter signal, a strong pump field $B_{\rm p}\cos(\omega_{\rm p}t)$ is applied parallel to $B_{\rm DM}$ with $B_{\rm p}\gg B_{\rm DM}$, in an approach similar to ruoso2016quax.
The magnetization density $\vec{m}(\vec{r},t)$ of the magnetized sample can be calculated using stationary solutions of the Bloch equations abragam1961principles,
$\frac{d\vec{m}}{dt}=\gamma\vec{m}\times\vec{B}-\frac{{m_{x}\hat{x}+m_{y}\hat{y}}}{T_{2}}-\frac{m_{z}-m_{0}}{T_{1}}\hat{z}$, (7)
as long as $|\omega_{\rm DM}-\omega_{\rm p}|\ll\omega_{L}$ augustine2002transient; ruoso2016quax. Here, $m_{0}=n_{s}\mu_{B}\tanh(\mu_{B}B_{0}/k_{B}T)$ is the equilibrium magnetization density, with $n_{s}$ as the spin number density, $T_{1}$ and $T_{2}$ as the longitudinal and transverse relaxation times. Due to the nonlinearity of the Bloch equations, the superposition of the two radio frequency fields, $B_{\rm p}cos(\omega_{\rm p})+B_{\rm DM}cos(\omega_{\rm DM})$, produces an on-resonance ($\gamma B_{0}-\omega_{\rm DM}=0$) beat-note modulation of $m_{z}$,
$\Delta m_{z}=m_{z}-m_{0}\simeq m_{0}\gamma^{2}T_{1}T_{2}B_{\rm p}B_{\rm DM}\cos(\omega_{\rm D}t),$ (8)
where $\omega_{\rm D}=|\omega_{\rm p}-\omega_{\rm DM}|$ being the detection frequency satisfying the condition $\omega_{\rm D}^{-1}<\min\left\{T_{1},T_{2}\right\}$. Expression 8 holds at resonance and for dark matter frequency around $\rm 1~GHz$ $(\omega_{L}=\omega_{\rm DM}=\rm GHz)$ where the dipole radiation damping effects, mentioned in ruoso2016quax, can be ignored. While a stronger pump field leads to larger amplification, one must ensure that the system is kept far from saturation $B_{\rm p}\ll 1/\gamma\sqrt{T_{1}T_{2}}$. For further details on the detection scheme, see Appendix B.
Choosing $\omega_{\rm D}=\omega_{m}$, this scenario becomes formally equivalent to applying an AM scheme in MRFM, with Eq. 6 giving the net force on the mechanical resonator,
$F_{z}(t)=m_{0}\gamma^{2}T_{1}T_{2}B_{\rm p}B_{\mathrm{\rm DM}}\cos(\omega_{\rm D}t)\int\frac{\partial B_{z}(\vec{r})}{\partial z}d^{3}r.$ (9)
The integral is performed over the resonant part of the sample volume, where the electron spins are in magnetic resonance as defined by the condition $\omega_{\rm DM}=\omega_{L}\pm\delta\omega$ with linewidth given by $\delta\omega=\frac{1}{T_{2}}$ far from saturation. Near the zero-probe-field resonance (ZPFR) regime, the entire sample can be within the resonance region suter2004magnetic, which gives a large force on the mechanical system, despite the weak field gradient. To determine the sensitivity of a MRFM setup to radio frequency field $B_{\rm DM}$, we now consider the contribution of various noise sources to the resonator motion.
Figure 2: Sensitivity plot of axion-electron coupling strength $g_{aee}$ as a function of axion mass $m_{\rm DM}$ (lower axis) and corresponding Compton frequency $f_{\rm DM}$ (top axis). A second y-axis gives the strength of the effective magnetic field induced on electrons by axions corresponding to the coupling strength $g_{aee}$. Our results for MRFM based axion search are given in red and blue for a membrane and cantilever, respectively. For each resonator, a 100 day long run at a single frequency is shown as well as a 100 day scan of minute long runs at neighboring frequencies, showing that one can choose to increase sensitivity or bandwidth. The QCD axion band is shown in yellow and existing constraints from astrophysical capozzi2020axion, space gondolo2009solar; aprile2022search and laboratory experiments terrano2015short; yan2019constraining are shown in gray. Also shown are narrowband constraints from other magnetic resonance based searches in green, labeled QUAX crescini2018operation; crescini2020axion, UWA flower2019broadening and QND ikeda2022axion. The top right panel shows the finer details of a single sensitivity line including the two sidebands at $\omega_{p}\pm\omega_{m}$. The bottom right panel shows the contributions from the back action noise (purple), thermal noise (yellow), and imprecision noise (green) to the total noise equivalent force (red).
Figure 3: Sensitivity plot of dark photon kinetic coupling parameter $\epsilon$ as a function of dark photon mass $m_{\rm DM}$ (lower axis) and corresponding Compton frequency $f_{\rm DM}$ (upper axis). Our results for MRFM based axion search are given in red and blue for a membrane and cantilever, respectively. For each resonator, a 100 day long run at a single frequency is shown as well as a 100 day scan of minute long runs at neighboring frequencies, showing that one can choose to increase sensitivity or bandwidth. The gray region represents constraints from astrophysical and cosmological observations assuming the dark photon makes up majority of dark matter, while dark green and light green are constraints from haloscope searches and rescaled limits from axion searches. See caputo2021dark for details of various constraints on the dark photon kinetic coupling.
Noise sources and minimum detectable coupling.– At mechanical resonance $\omega\sim\omega_{m}$, the sum of the measurement back action and imprecision noise is at a minimum, and the dominant noise source is due to the thermal bath of the mechanical system, with noise spectral density
$S_{FF}^{\rm th}=\frac{4k_{B}Tm_{\rm eff}\omega_{m}}{Q},$ (10)
where $T$, $Q$ and $m_{\rm eff}$ denote the temperature, quality factor and effective mass of the loaded mechanical oscillator, respectively. After an integration time of $\tau_{\rm int}$, the minimum detectable force is then given by
$F_{\rm noise}(\omega)=\frac{1}{(t_{\rm int}\tau_{\rm c})^{1/4}}\sqrt{S_{FF}^{\rm NEF}(\omega)},$ (11)
where the dark matter coherence time is $\tau_{c}\equiv\tau_{\rm DM}=\frac{c^{2}}{v_{\rm DM}^{2}f_{\rm DM}}=\frac{10^{6}}{f_{\rm DM}}$ for couplings proportional to the field and $\tau_{c}\equiv\tau_{\nabla a}=0.68\ \tau_{\rm DM}$ for couplings proportional to the gradient of the axion field $a$ barbieri2017searching. Here, NEF denotes the noise equivalent force, which corresponds to a signal to noise ratio of 1.
One can use Eq. 11 with Eq. 10 to calculate the thermal noise contribution to the force after integrating for $t_{\rm int}$. Assuming a signal-to-noise ratio of 1, we can then use Eq. 9 to determine the minimum detectable magnetic field using our proposed setup,
$B_{\rm DM}^{\rm sen}(\omega_{m})=\frac{\frac{1}{(t_{\rm int}\tau_{\rm c})^{1/4}}\sqrt{\frac{4k_{B}Tm_{\rm eff}\omega_{m}}{Q}}}{m_{0}\gamma^{2}T_{1}T_{2}B_{\rm p}\int{\frac{\partial B_{z}(\vec{r})}{\partial z}d^{3}r}}.$ (12)
The mechanical resonator has worse sensitivity away from the mechanical resonance due to large imprecision noise (see Appendix C for details). These expressions can be applied to any magnetically loaded mechanical system, provided that the underlying approximations hold.
Unlike several other magnetic mechanical systems, eddy-current effects do not pose a limitation here due to a combination of frequency mismatch and the material choice of YIG, which is not a conductor.
DM Parameter Space.– Before providing specific details of the device used for our estimates, we now overview existing constraints on the ULDM parameter space to work out the required magnetic field sensitivity. The parameter space for axion-electron coupling with existing laboratory and astrophysical constraints is shown in Fig. 2. The most stringent constraint on the axion-electron coupling at $g_{ae}\sim 10^{-13}~\rm eV$ stems from the lack of observed reduction in the tip of the red giant branch (TRGB) luminosity in globular clusters, where axion bremsstrahlung from electrons would otherwise cause significant energy loss and cooling capozzi2020axion. Weaker constraints arise from solar axion searches, including limits on axion-induced solar energy losses from measurements of the neutrino flux gondolo2009solar and null detection of electron recoils from solar neutrinos aprile2022search. The best current laboratory constraint in the relevant mass range is due to torsion pendulum measurements of dipole-dipole forces between electrons terrano2015short at $7\times 10^{-9}~\rm eV$, while measurement of the anomalous magnetic moment of electrons yan2019constraining also gives weaker constraints.
Three major milestones in the search for axions via their electron coupling are probing the coupling better than torsion pendulum limits terrano2015short, observations of tip of the red giant branch (TRGB) luminosity in globular clusters capozzi2020axion, and probing the QCD axion bands for DFSZ and KSVZ-like axions. These correspond to magnetic field strengths of $3\times 10^{-16}~\rm T$, $5\times 10^{-21}~\rm T$, and $2\times 10^{-24}~\rm T$, respectively (see Fig. 2). The best current narrowband experimental constraints are from magnetic axion haloscopes that rely on detecting the axion-magnon coupling crescini2018operation; crescini2020axion; flower2019broadening; ikeda2022axion. A number of proposals relying on optimistic advancements in technology have been proposed using QND measurements of axion-magnon conversion chigusa2020detecting; mitridate2020detectability; berlin2024physical; catinari2025hunting; chang2025mosaic. Our MRFM proposal for axion searches is similar in technical requirements and detection sensitivity to the best existing laboratory constraints, including QUAX crescini2018operation; crescini2020axion. Unlike the aforementioned proposals, it does not require major technological advancements and the experiment can be performed with current technology.
Device parameters.– As illustrated in Fig 1, displacement of the micromagnet-loaded mechanical oscillator with frequency $\omega_{m}$ is detected optically via an interferometry setup. The dark matter frequency $\omega_{\rm DM}$ searched for by the experiment can be tuned by varying $\omega_{\rm p}$, the pump field frequency. Our scheme allows for a scanning search for axions using the same mechanical system, as mechanical resonance frequencies cannot be tuned significantly in most systems.
Using the parameters given in Table 1 and equation 11, we find a minimum magnetic field sensitivity at resonance of about $2.0\times 10^{-16}$ T ($1.0\times 10^{-17}$ T) for a minute (100 days) of integration time with a membrane and $2.4\times 10^{-16}$ T ($1.2\times 10^{-17}$ T) for a minute (100 days) of integration time with a cantilever. Using Equations 1 and 4, we obtain minimum values that can be probed by our setup of the axion-electron coupling and the dark photon kinetic mixing parameters, respectively. These are plotted in Figs. 2 and 3. For the axion-electron coupling, an MRFM experiment can search for axions beyond existing laboratory constraints. For the dark photon kinetic mixing, this system allows for filling in gaps between existing haloscope constraints.
Parameters Used
Temperature $T$ 0.01K0.01~K
Laser Power $P$ $10^{-6}~\rm W$
Fabry-Perot Cavity
Resonant Frequency $\omega_{\rm cav}$ $1.77\times 10^{15}~\rm rad/s$
Length $L_{\rm cav}$ $10^{-4}~\rm m$
Finesse $F_{\rm cav}$ 200200
Fe-Co Micromagnet
Saturation Magnetization $M_{0}$ 2.4A/m2.4~A/m
Radius $R_{\rm mag}$ $100~\rm\mu m$
Density $\rho_{\rm mag}$ $8.5~\rm g/cm^{3}$
Mass $m_{\rm mag}$ $36~\rm\mu g$
SiN Membrane
Quality Factor $Q_{\rm m}$ $10^{6}$
Resonant Frequency $\omega_{\rm m}$ $2\pi\times 5.8~\rm kHz$
Cantilever
Quality Factor $Q_{\rm m}$ $10^{5}$
Resonant Frequency $\omega_{\rm m}$ $2\pi\times 1.1~\rm kHz$
YIG Sample
Longitudinal Relaxation Time $T_{1}$ $10^{-6}~\rm s$
Transverse Relaxation Time $T_{2}$ $10^{-6}~\rm s$
Number density $N_{\rm sample}$ $2.1\times 10^{28}~\rm m^{-3}$
Radius $r_{\rm sample}$ $1~\rm mm$
Table 1: Experimental parameters used in our calculations of the sensitivity of MRFM to dark matter-induced magnetic fields. The values for the resonators are chosen in the range of parameters from scozzaro2016magnetic; Takahashi2018; zwickl2012progress for membrane and wago1998paramagnetic for cantilever.
The bottom right panel of Fig. 2 shows contributions to the noise equivalent force spectral density from back action noise (purple), thermal noise (yellow) and imprecision noise (green). The top right panel shows the corresponding plot for sensitivity to the axion-electron coupling, showing zoomed in part of both sidebands $\omega_{p}\pm\omega_{m}$. This two-sideband feature comes from the frequency up conversion scheme and has been seen by, for example, the AURIGA detector branca2017search. Our proposed experiment is limited by thermal noise over a relatively broad band with a linewidth of around $100~\rm Hz$ (see Fig. 5 in Appendix C).
The target dark matter frequency $\omega_{\rm DM}$ can be scanned by varying the pump frequency $\omega_{\rm p}$ and the bias field $B_{0}$. Since the magnetic resonance linewidth ($\sim\rm MHz$) is much larger than the mechanical frequency, a significant scanning region can be probed with the same bias field value. With minute long runs, one can scan a $14.4~\rm MHz$ frequency range in $100$ days.
One of the limits of the setup is the size of the sample. There is a limit to how large a YIG sphere can be fabricated, and large spheres may be too sensitive to adverse effects, such as radiation damping. Since our proposed setup allows for a large sample area, which translates to increased sensitivity to lower magnetic field values (see Eq. 12), one can use multiple YIG spheres, as done by QUAX crescini2020axion. A potential problem here may be the transfer of photons from radiation damping from one sphere to its neighbor. However, this problem can be resolved by insulating the spheres. In our calculations, we take a configuration of 9 spheres (see Fig 4). The number of YIG spheres, i.e. the number of spins, is not a technical limitation of the proposal, but will be limited for any experiment by available resources.
The constraints presented here involve just a minute of integration time on existing technology, with the entire experimental setup shown in figure 1 taking $\sim$1 cm3 of volume. It is possible to fit multiple such setups inside a dilution refrigerator, improving the sensitivity by $\sqrt{N_{e}}$, for $N_{e}$ experimental setups, or scanning over a much larger bandwidth. A higher mechanical quality factor, as demonstrated on several platforms, can allow one to probe even weaker ULDM couplings. There is also potential for further multiplexing by using more magnetized samples, which only requires one to adjust the applied bias field to widen the resonance region. Furthermore, for axion-electron couplings in particular, flux concentrators can be used to amplify the signal gao2025focusing. Finally, our proposal involves standard quantum limit readout, a limit that has already been circumvented in several experimental platforms, e.g. xia2023entanglement; jia2024squeezing; mason2019continuous.
Conclusion and outlook. – While most MRFM research has pushed towards detecting single spin flips, our work motivates exploration of another regime, namely MRFM as an ultra-weak magnetometer. This requires using a frequency scanning technique that can be adapted to other optomechanical searches for dark matter manley2020searching; manley2021searching; hirschel2024superfluid. Our proposal for a magnetic resonance force microscope search for dark matter induced magnetic fields is just one thrust within the larger optomechanics for fundamental physics theme carney2021mechanical; moore2021searching; kalia2024ultralight; ahrens2025levitated; bose2025massive. Beyond dark matter, other interesting signals from gravitational waves, neutrinos, etc. can potentially be probed using optomechanical systems, paving the way for pushing forward technology and potentially discovering new physics.
I Acknowledgments
This work is supported by the National Science Foundation grant PHY-2047707 and the Office of the Under Secretary of Defense for Research and Engineering under Award No. FA9550-22-1-0323. We thank Kylee Smith for preparing Fig. 1. We are also grateful to Yevgeny Stadnik and Jack Sankey for helpful discussions.
Appendix A Appendix A: Dark Matter Couplings
Axion-Fermion Coupling.– The axion field can be written as a classical wave in the form
$a(\vec{x},t)=\frac{\sqrt{2\rho_{\rm DM}}}{\omega_{\rm a}}\cos(\omega_{a}t+\vec{k}_{a}\cdot\vec{x}+\dots),$ (13)
where the dark matter Compton frequency is related to its mass as $\omega_{a}=\frac{m_{a}c^{2}}{\hbar}$ and $|\vec{k}_{a}|=\frac{m_{a}v_{a}}{\hbar}=\frac{10^{-3}m_{a}c}{\hbar}$. Axions can couple to fermions via the operator
$\mathcal{L}_{af}=\frac{g_{af}\sqrt{\hbar^{3}c}}{2m_{f}}\partial_{\mu}a\bar{\psi}_{f}\gamma^{\mu}\gamma_{5}\psi_{f}.$ (14)
Here $f=e,p,n$ for electrons, protons and neutrons. We will ignore nuclear structure when dealing with nuclei. This leads to an interaction Hamiltonian analogous to that of a magnetic field interacting with the fermion spin
$H_{af}=-\gamma_{f}\vec{S}_{f}\cdot\vec{B}_{af},$ (15)
where the gyromagnetic ratio is $\gamma_{f}=\frac{g_{f}e}{2m_{f}}$, the spin operator is $\vec{S}_{f}=\frac{\hbar}{2}\vec{\sigma}_{f}$ and the effective magnetic field is
$\vec{B}_{af}=\frac{2g_{af}\sqrt{\hbar c}}{g_{f}e}\frac{\vec{v}_{\rm DM}}{c^{2}}\sqrt{2\rho_{\rm DM}}\sin(\omega_{a}t),$ (16)
where the fermion g-factors are $g_{e}\approx 2,g_{\rm p}\approx 5.58,g_{n}\approx-3.8$ and we ignore the spatial variation of the axion field.
Dark Photon Kinetic Mixing.– The dark photon ${A^{\prime}}^{\mu}$ can kinetically mix with with the Standard Model photon $A^{\mu}$, leading to the following interaction operator in the interaction basis fedderke2021earth; kalia2024ultralight
$\mathcal{L}_{A^{\prime}}\supset\frac{\epsilon}{\mu_{0}}\left(\frac{m_{A^{\prime}}c}{\hbar}\right)^{2}A_{\mu}{A^{\prime}}^{\mu}=-J_{\rm eff}^{\mu}A_{\mu},$ (17)
where $\epsilon$ is the kinetic mixing parameter. The interaction is equivalent to one between the EM field $A_{\mu}$ and an effective current $J^{\mu}_{\rm eff}$. Applying the Euler-Lagrange equations with respect to $A^{\mu}$ to $\mathcal{L}\supset-\frac{1}{4}F^{\mu\nu}F_{\mu\nu}-J^{\mu}_{\rm eff}A_{\mu}$, we obtain the equation of motion $\partial_{\nu}F^{\mu\nu}=J^{\mu}_{\rm eff}$. Using the equation of motion for massive vector fields $\partial_{\mu}{A^{\prime}}^{\mu}=0$, taking the non-relativistic limit, one can see ${A^{\prime}}^{0}=0$. Then, the dark photon field can be written as higgins2024maglev
$|\vec{A}^{\prime}(\vec{x},t)|\approx\frac{\sqrt{2\rho_{\rm DM}}}{\omega_{A^{\prime}}}\cos(\omega_{A^{\prime}}t).$ (18)
Hence, the equation of motion $\partial_{\nu}F^{\mu\nu}=J^{\mu}_{\rm eff}$ gives the Ampere-Maxwell law
$\vec{\nabla}\times\vec{B}+\frac{1}{c^{2}}\frac{\partial\vec{E}}{\partial t}=\mu_{0}\vec{J}_{\rm eff}.$ (19)
Therefore, the dark photon induced effective current generates real electromagnetic fields. When the dark photon Compton wavelength $\lambda_{A^{\prime}}=c/f_{A^{\prime}}=2\pi\hbar/m_{A^{\prime}}c$ is much larger than the experiment size $L$ (i.e. size of magnetic shield), the spatial variation of $\vec{J}_{\rm eff}$ can be ignored. In this regime, the magnetic field part dominates over the electric field term, since
$\left|\frac{\partial_{t}\vec{E}/c^{2}}{\vec{\nabla}\times\vec{B}}\right|\sim\frac{1}{c^{2}}\frac{\omega_{A^{\prime}}|\vec{E}|}{|\vec{B}|/L}=\frac{2\pi L}{\lambda_{A^{\prime}}}\ll 1.$ (20)
Hence, the primary effect of the dark photon is an oscillating magnetic field of amplitude
$B_{A^{\prime}}\sim\mu_{0}J_{\rm eff}L\sim\frac{\sqrt{2\mu_{0}\rho_{\rm DM}}c}{\hbar}\epsilon m_{A^{\prime}}L.$ (21)
Axion-Photon Coupling.– Axions can also couple to photons via the following operator which can also be written as an interaction between the EM field and an effective current kalia2024ultralight
$\mathcal{L}_{a\gamma}=\frac{g_{a\gamma}\sqrt{\hbar c^{3}}}{4\mu_{0}}aF_{\mu\nu}\tilde{F}^{\mu\nu}=J^{\mu}_{\rm eff}A_{\mu},$ (22)
where $\tilde{F}^{\mu\nu}=\frac{1}{2}\epsilon^{\mu\nu\rho\sigma}F_{\rho\sigma}$ and the $J^{\mu}_{\rm eff}=-\frac{g_{a\gamma}\sqrt{\hbar c^{3}}}{\mu_{0}}\partial_{\nu}a\tilde{F}^{\mu\nu}$. Again, the zeroth component gives 0 and the current vector induces a magnetic field
$B_{a\gamma}=\sqrt{2\hbar c\rho_{\rm DM}}g_{a\gamma}LB_{0}.$ (23)
Note that unlike the dark photon, the axion-photon induced magnetic field requires a background mediating magnetic field $B_{0}$.
Appendix B Appendix B: MRFM Details
Comparing MRFM to magnetic resonance imaging.– Here we expand on some differences between magnetic resonance imaging and the MRFM scheme proposed here.
We first expand on inductive vs. mechanical detection. In conventional inductive detection, the precession of a sample’s magnetization induces a current in a pickup coil, which in turn generates a magnetic field that reacts back on the sample. This exchange of energy resulting in back-action noise, historically known as radiation damping, effectively damps the magnetization precession within the characteristic time $T_{R}=(2\pi\eta Q_{\rm coil}\gamma\mu_{0}M_{0})^{-1}$, where $\eta$ is the filling factor (volume of sample per volume of coil ($V_{s}/V_{c}$)) and $Q_{\rm coil}$ is the coil’s quality factor muller2007nuclear; ruoso2016quax. MRFM overcomes this limitation by interacting with the sample through a magnetic force gradient (a one-way force detection) rather than a reciprocal electromagnetic induction (a two-way energy exchange). This allows for a less perturbative and therefore more sensitive measurement with a better signal-to-noise ratio sidles1993signal.
While various modulation schemes enable detection of longitudinal magnetization, transverse magnetization detection in MRFM is challenging. This is due to the frequency mismatch between typical Larmor and mechanical frequencies. Even in the absence of this mismatch, spin control is complicated due to the requirement of a strong radio frequency field and non-applicability of the rotating wave approximation fischer2019spin.
Micromagnet.– Here we expand on the theoretical treatment of the micromagnet. An integral part of an MRFM setup is the micromagnet that generates an inhomogeneous magnetic field with a spatial gradient that couples the mechanical resonator to the sample magnetic moments. In our setup, the micromagnet is approximated as a micro-sphere with radius $R_{s}$ and has saturation magnetization $M_{0}$, uniformly magnetized along $\hat{z}$. The criterion for the selection of the radius $R_{s}$ is its ratio with the distance between the center of the micromagnet and the surface of the sample ($d$), such that $R_{s}/d\ll 1$, and hence the magnet creates a dipolar magnetic field with strength $B(r)=\frac{4\pi}{3}M_{0}(R_{s}/r)^{3}$ at radial distance $r$ from its center suter2002probe; hammel2007magnetic; sidles91. This gives the following field gradient
$\frac{\partial B(\vec{r})}{\partial z}$ $=\frac{-3B(r)}{2r}(3\cos(2\theta)+1)\cos(\theta)$ (24)
$+\frac{3B(r)}{r}\sin(2\theta)\sin(\theta),$
with the polar angle $\theta$ between the positive z-axis and the radial vector.
For the material of the micromagnet, we considered Fe-Co because of its high saturation magnetization $M_{0}\sim 2.4\quad A/m$ xue2011geometry; liu2008magnetic.
Micromagnet Loaded Mechanical System.– Here we discuss how to calculate the resonance frequency of a micromagnet loaded mechanical resonator. For example, consider a $V_{m}=\rm 5~mm\times 5~mm\times 500~nm$ volume SiN membrane ($\rho_{m}=3.17~\rm g/cm^{3}$) with spring constant $k=60~\rm N/m$ loaded with a micromagnet of mass $m_{\rm magnet}=36~\rm\mu g$. Then the mechanical resonator frequency can be calculated using $\omega_{m}=\sqrt{k/m_{\rm eff}}$, where the effective mass is approximately $m_{\rm eff}\approx m_{\rm magnet}+\frac{m_{\rm mech}}{4}$. Using the above numbers, the magnet-loaded membrane frequency is $\omega_{m}\approx 2\pi\times 5.8~\rm kHz$. For cantilever, we use the magnetically loaded resonance frequency $\omega_{m}=2\pi\times 1.1~\rm kHz$, which matches that of wago1998paramagnetic. In general, the resonance frequency of the loaded mechanical system will be estimated numerically.
Resonance Volume.– The micromagnet and the bias field $B_{0}$ define the region of space where magnetic resonance can be induced. For conditions $B_{0}\neq B_{\rm rf}/\gamma$, the sensitive slice is a thin shell with spatial thickness $x_{sl}\approx\frac{\delta B}{\nabla B}$, where $\delta B\propto\frac{1}{T_{2}}$ is the intrinsic resonance linewidth of the sample. However, this concept breaks down when $B_{0}=B_{\rm rf}/\gamma$, where the influence of the gradient becomes small across most of the sample, hence known as zero-probe-field resonance (ZPFR). In this regime, the sensitive region is no longer a slice and expands to cover nearly the entire sample. Thus, each spin experiences a total field close to $B_{0}$, and this results in a strong signal despite the weak local field suter2002probe.
Changes in longitudinal magnetization.– The evolution of an ensemble of spins in a magnetized sample due to external magnetic fields (magnetizing field $B_{0}\hat{z}$, and a driving radio-frequency field $B_{\rm rf}$ in the x-y plane) is described by Bloch equations abragam1961principles:
$\frac{d\vec{m}}{dt}=\gamma\vec{m}\times\vec{B}-\frac{{m_{x}\hat{x}+m_{y}\hat{y}}}{T_{2}}-\frac{m_{z}-m_{0}}{T_{1}}\hat{z}$, (25)
where $\vec{m}$ is the magnetization density vector, $\vec{m}=(m_{x},m_{y},m_{z})$.
For low d.c. fields, comparable in magnitude to the r.f. field, one instead considers the modified Bloch equations $\frac{d\vec{m}}{dt}=\gamma\vec{m}\times\vec{B}-\frac{m_{z}-\chi_{0}(B_{0}+B_{z})}{T_{1}}\hat{z}-\frac{m_{x}\hat{x}+m_{y}\hat{y}}{T_{2}}$ abragam1961principles. The longitudinal steady-state solution is
$m_{z}=\frac{\chi_{0}}{\mu_{0}}B_{0}\big(1+\frac{(\omega/B_{0}\gamma)(\gamma B_{\rm rf})^{2}T_{1}T_{2}}{1+(T_{2}\Delta\omega)^{2}+(\gamma B_{\rm rf})^{2}T_{1}T_{2}}\big),$ (26)
where $\frac{\chi_{0}}{\mu_{0}}B_{0}=m_{0}$, $B_{0}\gamma=\omega_{L}$, $\Delta\omega=\omega_{L}-\omega_{\rm rf}$. At low driving field (as defined by $\gamma^{2}B_{\rm rf}^{2}T_{1}T_{2}\ll 1$), the spins excited at resonance ($\omega=\omega_{L}=\omega_{\rm rf}$) vary the longitudinal magnetization to $m_{0}+\Delta m_{z}$, where $\Delta m_{z}=m_{0}\gamma^{2}B_{\rm rf}^{2}T_{1}T_{2}$ wago1997magnetic.
In our scenario ($B_{\rm rf}=B_{\rm p}+B_{\rm DM}$), leading to
$B_{\rm rf}^{2}$ $=\frac{B_{\rm p}^{2}}{2}[1+\cos(2\omega_{\rm p}t)]+\frac{B_{\rm DM}^{2}}{2}[1+\cos(2\omega_{\rm DM}t)]$
$+2B_{\rm p}B_{\rm DM}\frac{\cos((\omega_{\rm p}-\omega_{\rm DM})t)+\cos((\omega_{\rm p}+\omega_{\rm DM})t)}{2}.$ (27)
Here, The D.C. terms $B_{\rm p}^{2}$ and $B_{\rm DM}^{2}$ give constant shifts, which are not useful for detection. Meanwhile, the high-frequency terms $\cos(2\omega_{\rm p}t)$, $\cos(2\omega_{\rm DM}t)$ and $\cos((\omega_{\rm p}+\omega_{\rm DM})t)$, oscillating at frequencies of the order of GHz, are inaccessible to the mechanical resonator. Therefore, only the difference frequency term $\cos((\omega_{\rm p}-\omega_{\rm DM})t)$ contributes to the signal. Since $\omega_{\rm p}\approx\omega_{\rm DM}$ and $|\omega_{\rm p}-\omega_{\rm DM}|\ll\omega_{L}$, the difference $\omega_{\rm D}=|\omega_{\rm p}-\omega_{\rm DM}|$ is small and can be detected by a mechanical resonator if the $\omega_{\rm D}$ coincides with the mechanical resonance $\omega_{m}=\omega_{\rm D}$. The $\omega_{\rm D}$ frequency component of $\Delta m_{z}$ thus obtained is given in equation 7 of the main text.
Magnetic-Dipole radiation damping.- For a magnetic sample in free space with a high Larmor frequency and large magnetization $M_{0}$, the dominant loss mechanism is the magnetic dipole emission. This is an intrinsic quantum phenomenon where the precessing magnetization spontaneously emits electromagnetic radiation characterized by the relaxation time $\tau_{r}=\frac{\lambda_{L}^{3}}{\gamma\mu_{0}M_{0}V_{s}}$, where $\lambda_{L}=c/\nu_{L}$ is the Larmor wavelength abragam1961principles; crescini2019towards; ruoso2016quax. To avoid faster than spin-spin relaxation, the spherical YIG sample with Larmor frequency $1~\rm GHz$ ($10~\rm GHz$) must be limited in size to a maximum radius of around $9~\rm mm$ ($0.6~\rm mm$). Consequently, we restrict our investigation to frequencies around a $\rm GHz$ to avoid this coherence loss channel. Alternatively, strategies like embedding the sample in a resonant cavity (as done by QUAX) may be employed to manage the phase space of the emitted radiation crescini2019towards; ruoso2016quax; crescini2018operation.
Eddy-current effects.- A possible concern in MRFM experiments is the noise from eddy-current effects in either the sample or the micromagnet, especially in A.M. methods. In our proposed setup, however, this effect can be neglected for two reasons. Firstly, a YIG sample has low conductivity and, therefore, negligible eddy currents. Secondly, the applied pump field oscillates at $\rm GHz$ frequencies, but the mechanical resonator operates in the $\rm kHz$ band, creating a large mismatch between the frequency of the induced eddy fields by the pump and the frequency for force detection via the resonator. As a result, even with a stronger pump field, the fast oscillating eddy fields generated inside the micromagnet remain outside the mechanical detection bandwidth.
Figure 4: The magnetic field gradient produced by the resonator-attached micromagnet (shown as a red dot at the origin) is plotted as a contour against distance. The yellow region denotes the sensitive slice that is determined by the magnetic resonance condition $B_{0}+B_{z}=\frac{\omega_{L}(\vec{r})}{\gamma}$ and the linewidth $1/T_{2}$. The red circles depict a $3\times 3$ grid of YIG spheres.
Appendix C Appendix C: Force Equivalent Noise
The transfer function between force and position is obtained by solving the equations of motion for the system in the frequency domain aspelmeyer2014cavity
$\chi_{xF}(\omega)=\frac{x(\omega)}{F_{\rm ext}}=\frac{1/m}{(\omega_{m}^{2}-\omega^{2})+i\frac{\omega\omega_{m}}{Q}}.$ (28)
The oscillator response is given by
$S_{xx}(\omega)$ $=S_{xx}^{\rm imp}(\omega)+|\chi_{xF}(\omega)|^{2}\times$
$\left[S_{FF}^{\rm th}(\omega)+S_{FF}^{\rm ba}(\omega)+S_{FF}^{\rm\,signal}(\omega)\right].$ (29)
The noises in the system are
$S^{\rm imp}_{xx}(\omega)=2\frac{\kappa}{16\bar{n}G^{2}}\left(1+\left(\frac{\omega}{\kappa/2}\right)^{2}\right),$ (30)
$S_{FF}^{\rm ba}(\omega)=8\bar{n}\frac{4\hbar^{2}G^{2}}{\kappa}\frac{(\kappa/2)^{2}}{2}\frac{2}{(\kappa/2)^{2}+\omega^{2}},$ (31)
and
$S_{FF}^{\rm th}(\omega)=\frac{4k_{B}T}{\omega}\mathcal{I}m\left(\frac{1}{\chi_{xF}(\omega)}\right)$ (32)
saulson1990thermal, which are the imprecision, backaction, and thermal noise respectively.
Here, $G=\omega_{c}/L$ is the optomechanical coupling aspelmeyer2014cavity with cavity frequency $\omega_{c}$ and cantilever length $L$, $\kappa=\frac{\omega_{\rm FSR}}{F}=\frac{c}{2LF}$ with finesse $F$ is the cavity decay rate, and
$\bar{n}=\frac{\kappa P}{\hbar\omega_{c}(\kappa^{2}/4)},$ (33)
is the mean occupation of the cavity, where $P$ is the laser power.
The imprecision noise is a shot noise from the discrete nature of photons, the backaction noise is due to the radiation pressure of the laser, and the thermal noise is due to thermal fluctuations of the system. While not obvious from equation 32, the thermal noise in the system is white.
We now define a quantity known as the noise equivalent force (NEF). This is a force signal power such that it doubles the power at every frequency compared to the case of zero signal. In essence, this is a signal-to-noise-ratio (SNR) equal to 1 condition on the signal. It tells us how strong a possible signal would have to be at any frequency to be considered detectable. It can be found to be
$S_{FF}^{\rm NEF}(\omega)=|\chi_{xF}(\omega)|^{-2}S_{xx}^{\rm imp}(\omega)+S_{FF}^{\rm th}(\omega)+S_{FF}^{\rm ba}(\omega).$ (34)
The smallest force amplitude detectable in integration time $t_{\rm int}$ is given by
$F_{\min}(\omega)=\frac{1}{(t_{\rm int}\tau_{\rm c})^{1/4}}\sqrt{S_{FF}^{\rm NEF}(\omega)}.$ (35)
The quarter-root scaling with integration time is due to the fact that the coherence time of the signal is much less than the integration time. Through a method of averaging the power spectral densities (PSDs) of several short-time measurements of length equal to 1 coherence time (known as Bartlett’s method proakis2006), we arrive at the quarter-root scaling manley2023searching.
Figure 5: Noise contributions to the mechanical resonator motion due to back action noise (yellow), thermal noise (red), and imprecision noise (blue) are shown as the square root of the noise equivalent force spectral density as a function of frequency. The dominant part of the noise equivalent force (NEF), shown in black, is imprecision noise off-resonance and thermal noise on-resonance with the mechanical frequency $f_{m}$. The box on the lower left shows the zoomed in region around $f_{m}$. The bandwidth of the setup can be seen to be around $100~\rm Hz$.